### v1

@@ 7 Conclusion
In this study, we tackle the challenge of denoising task difficulty within the diffusion model framework and introduce a novel task difficulty-based curriculum learning approach. To the best of our knowledge, we are the first to define task difficulty by considering both the convergence rates of loss and performance metrics. Moreover, in terms of data distribution analysis, we observe a reduction in relative entropy between consecutive probability distributions as timesteps progress. We believe that these observations might help reorganize the conflicts of previous works regarding denoising task difficulties. Building upon these insights, we propose a curriculum learning framework for the diffusion model, comprising curriculum design and pacing strategies. Our experimental results convincingly demonstrate the efficacy of our approach across diverse diffusion model designs, datasets, and tasks. From these results, we emphasize that considering an order of learning denoising tasks is also a potential direction to improve the training of diffusion models. In future research, for further enhancements, more advanced curriculum learning strategies such as self-pacing can be elaborated.
A. Extended Related Work
@@
B. Detailed Experimental Setups for Observation
In Section 4, we examined the difficulty of denoising tasks in terms of convergence with various models $\{\mathrm{M}\}_{i=1}^{20}$, which are trained within specific timesteps $[\frac{i-1}{20}T,\frac{i}{20}T]$. For the DiT architecture, we employed the DiT-B/2, whereas for EDM, we used the DDPM++ architecture. Both DiT and EDM models were trained on the FFHQ dataset, with a batch size of 256, for approximately 20,000 and 4,000 iterations (equivalent to processing 1 million images), respectively. This training was conducted until both loss and performance converged.
To evaluate the performance of diffusion models through the FID score of the generated images, performing a recursive denoising task from $T$ to zero is necessary, complicating the assessment using only $\mathrm{M}_{i}$. Following [11], we generated samples where $\mathrm{M}_{i}$ was specifically utilized for denoising within its trained intervals. At the same time, the diffusion model is responsible for the denoising tasks across the entire range of timesteps. For this evaluation, we sampled 10,000 images using a DDIM sampler over 250 steps for DiT, and an Euler solver over 40 steps for the other models.
@@ B. Detailed Experimental Setups for Observation
Fig. A: Qualitative comparison between vanilla, naive curriculum, and ours on FFHQ dataset.
@@
Fig. B: Qualitative comparison between vanilla, naive curriculum, and ours on ImageNet dataset.
C. Experiment Details
For training diffusion models, we utilized the AdamW optimizer [36] with a constant learning rate of 0.0001, and weight decay was not applied. The exponential moving average (EMA) to the model’s weights was used to stabilize the training and the decay ratio was set to 0.9999. The batch size was set to 256, and we augmented the training data by a horizontal flip. While the diffusion timestep $T$ was configured as 1,000 for all experiments, we trained for 100K iterations for the FFHQ dataset [23], and 400K iterations for the ImageNet dataset [7] and MS-COCO dataset [33]. The number of clusters $N$ was 20 unless otherwise specified. The maximum patience $\tau$ was varied across model sizes: it was set at 200 for DiT-S/2, DiT-B/2, and EDM, and 400 for DiT-L/2. EDM was trained using fp16, while the other models were trained using fp32.
Unless otherwise mentioned, we sampled 50K samples for evaluation. In tasks involving conditional generation, including class-conditional image generation (e.g. CIFAR-10, ImageNet) and text-to-image conversion (e.g. MS-COCO), we adapted the classifier-free guidance [19] with a guidance scale of 1.5.
@@ C. Experiment Details
“A man is on a path riding a horse.”
“Tower of a brick building with a clock on the front.”
“A dog leaps through the air as it catches a frisbee.”
“A person in white and red snow suit skiing on a slope.”
“A group of young people playing a game of soccer.”
“A young boy is smiling and he has food around him on a table.”
“a small boy with a hat is standing on a surf board.”
Fig. C: Qualitative comparison between vanilla, naive curriculum, and ours on MS-COCO dataset.
D. Qualitative Results
7.0.1 Qualitative Evaluation on the FFHQ Dataset.
@@ D. Qualitative Results
7.0.2 Qualitative Analysis on the ImageNet Dataset.
@@
7.0.3 Qualitative Assessment on the MS-COCO Dataset.
@@
E. Further Experimental Results
E.1. Convergence Speed across Model Size
Fig. D: We observed an increase in convergence speed across various model sizes when the proposed curriculum learning approach was applied.
@@ E.1. Convergence Speed across Model Size
E.2. Robustness on Noise Schedule
@@
For more comprehensive ablation study, we also trained the diffusion model with different noise schedule. In contrast to cosine scheduling, the $\beta_{t}$ is set by uniformly dividing the interval $[0.0001,0.02]$, and the $C_{i}$ are obtained corresponding to SNR on a linear schedule. As shown in Table. A, our approach improves the performance with cosine and linear noise schedulers both.
F. Negative Social Impacts

### v2

@@ 7 Conclusion
In this study, we tackle the challenge of denoising task difficulty within the diffusion model framework and introduce a novel task difficulty-based curriculum learning approach. To the best of our knowledge, we are the first to define task difficulty by considering both the convergence rates of loss and performance metrics. Moreover, in terms of data distribution analysis, we observe a reduction in relative entropy between consecutive probability distributions as timesteps progress. We believe that these observations might help reorganize the conflicts of previous works regarding denoising task difficulties. Building upon these insights, we propose a curriculum learning framework for diffusion models, comprising curriculum design and pacing strategies. Our experimental results convincingly demonstrate the efficacy of our approach across diverse diffusion model designs, datasets, and tasks. From these results, we emphasize that considering an order of learning denoising tasks is also a potential direction to improve training of diffusion models. In future research, for further enhancements, more advanced curriculum learning strategies such as self-pacing can be elaborated.
Appendix
Appendix A Extended Related Work
@@
Appendix B Detailed Experimental Setups for Observation
In Section 4, we examined the difficulty of denoising tasks in terms of convergence with various models $\{\mathrm{M}\}_{i=1}^{20}$, which are trained within specific timesteps $[\frac{i-1}{20}T,\frac{i}{20}T]$ for DiT and SiT, and $[\mathrm{\Phi}^{-1}(\frac{i-1}{N}),\mathrm{\Phi}^{-1}(\frac{i}{N})]$ for EDM where $\mathrm{\Phi}^{-1}$ is the inverse cumulative distribution function of the Gaussian distribution. For the DiT architecture, we employed the DiT-B/2, whereas for EDM, we used the DDPM++ architecture. Both DiT and EDM models were trained on the FFHQ dataset, with a batch size of 256, for approximately 20,000 and 4,000 iterations (equivalent to processing 1 million images), respectively. This training was conducted until both loss and performance converged. Configuration of training optimizers and learning rates are the same as setups in Section 6.
To evaluate the performance of diffusion models through the FID score of the generated images, performing a recursive denoising task from $T$ to zero is necessary, complicating the assessment using only $\mathrm{M}_{i}$. Following [11], we generated samples where $\mathrm{M}_{i}$ was specifically utilized for denoising within its trained intervals. At the same time, the diffusion model is responsible for the denoising tasks across the entire range of timesteps. For this evaluation, we sampled 10K images using a DDPM sampler over 250 steps for DiT and SiT, and an Euler solver over 40 steps for the other models.
@@ Appendix B Detailed Experimental Setups for Observation
Figure A: Qualitative comparison between vanilla, naive curriculum, and ours on the FFHQ dataset.
@@
Figure B: Qualitative comparison between vanilla, naive curriculum, and ours on ImageNet dataset.
Appendix C Approximation of KL Divergence of $p_{t-1}$ and $p_{t}$.
Here, we supplement the approximation of KL Divergence of $p_{t-1}$ and $p_{t}$ omitted in Section 4.2. To explore the difficulties of denoising tasks from the distributional viewpoint, we analyze the KL divergence of $p_{t-1}$ and $p_{t}$, $D_{KL}(p_{t-1}||p_{t})$. However, due to the unknown explicit density form of $p_{0}$, it is approximated through unbiased estimators as follows:
where $\hat{D}_{KL}$ and $\hat{p}_{t}$ are unbiased estimators of $D_{KL}$ and $p_{t}$, respectively, and we choose Monte-Carlo estimators for them [10, 39].
Appendix D Algorithm
Due to the limited space of the main manuscript, we hereby present the step-by-step process of our method to supplement the details of our approach. The pacing function, which determines the moments to transit between curriculum stages is described in Algorithm 1. By incorporating this pacing function, the detailed procedure of our proposed curriculum learning method for training diffusion is illustrated in Algorithm 2.
Algorithm 1 Pacing Function
Input: Current loss $L_{\text{cur}}$, Best loss $L_{\text{best}}$, Current patience $\tau_{\text{cur}}$, Maximum patience $\tau_{\text{max}}$, Current curriculum index $I_{\text{cur}}$
Output: Updated patience, Updated curriculum index
# Reset patience
if $L_{\text{cur}}<L_{\text{best}}$ then
    return 0, $I_{\text{cur}}$
else
    # Proceed to next curriculum
    if $\tau_{\text{cur}}+1>\tau_{\text{max}}$ then
       return 0, $I_{\text{cur}}-1$
   # Increase patience
    else
       return $p_{\text{cur}}+1,I_{\text{cur}}$
    end if
end if
Algorithm 2 Curriculum Learning
Input: Curriculum $\{C_{i}\}_{i=1}^{N}$, Pacing function $g$, Maximum patience $\tau_{\text{max}}$, Loss function $f$, Curriculum index $I_{\text{cur}}=N$, Best loss $L_{\text{best}}=\infty$, Model $M_{\theta}$
while $I_{\text{cur}}>0$ do
    # Mini-batch sampling
    $X\sim C_{I_{\text{cur}}}$
    # Calculate Loss
    $L_{\text{cur}}=f(M_{\theta}(X))$
    # Update model
    $\theta$ = $\theta$ $\nabla_{\theta}{L_{\text{cur}}}$
    # Pacing function
    $\tau_{\text{cur}},I_{\text{next}}=g(L_{\text{cur}},L_{\text{best}},\tau_{\text{cur}},\tau_{\text{max}},I_{\text{cur}})$
    # Update curriculum
    if $I_{\text{cur}}\neq I_{\text{next}}$ then
       $I_{\text{cur}}=I_{\text{next}}$
       $L_{\text{best}}=\infty$
   # Update best loss
    else if $L_{\text{cur}}<L_{\text{best}}$ then
       $L_{\text{best}}=L_{\text{cur}}$
    end if
end while
Appendix E Details on Experimental Setups
Evaluation metrics.
To evaluate the performance of models, we utilized three metrics: FID [16], IS [51], and Precision/Recall [30]. Specifically, we applied FID and IS to measure sample quality, while Precision is used to assess quality further and Recall was utilized to evaluate the diversity of the generated samples in ImageNet setup. In other datasets, we employed FID to evaluate sample quality. Unless otherwise mentioned, we sampled 50K samples for evaluation. In tasks involving conditional generation, including class-conditional image generation (e.g. CIFAR-10, ImageNet) and text-to-image conversion (e.g. MS-COCO), we adapted the classifier-free guidance [19] with a guidance scale of 1.5.
Training details.
For training diffusion models, we utilized the AdamW optimizer [36] with a constant learning rate of 0.0001, and weight decay was not applied. The exponential moving average (EMA) to the model’s weights was used to stabilize the training and the decay ratio was set to 0.9999. The batch size was set to 256, and we augmented the training data by a horizontal flip. While the diffusion timestep $T$ was configured as 1,000 for all experiments, we trained for 100K iterations for the FFHQ dataset [24], and 400K iterations for the ImageNet dataset [7] and MS-COCO dataset [33]. The number of clusters $N$ was 20 unless otherwise specified. The maximum patience $\tau$ was varied across model sizes: it was set at 200 for DiT-S/2, DiT-B/2, and EDM, and 400 for DiT-L/2. EDM was trained using fp16, while the other models were trained using fp32. We used 8 A100 GPUs for all experiments.
@@ Appendix E Details on Experimental Setups
“A man is on a path riding a horse.” “Tower of a brick building with a clock on the front.” “A dog leaps through the air as it catches a frisbee.” “A person in white and red snow suit skiing on a slope.” “A group of young people playing a game of soccer.” “A young boy is smiling and he has food around him on a table.” “a small boy with a hat is standing on a surf board.”
Figure C: Qualitative comparison between vanilla, naive curriculum, and ours on MS-COCO dataset.
Appendix F Qualitative Results
In this section, we present qualitative comparisons between three methods: 1) Vanilla, 2) NaiveCL, and 3) Ours, across the FFHQ, ImageNet, and MS-COCO datasets. All methods are evaluated using DiT-B models, and the final trained models generate all samples shown in the results. As shown in the results in the following subsections, our approach can synthesize more accurate and realistic images compared to Vanilla and NaiveCL.
F.1 Qualitative Evaluation on the FFHQ Dataset.
@@ Appendix F Qualitative Results
F.2 Qualitative Analysis on the ImageNet Dataset.
@@
F.3 Qualitative Assessment on the MS-COCO Dataset.
@@
Appendix G Further Experimental Results
G.1 Convergence Speed across Model Size
Figure D: We observed an increase in convergence speed across various model sizes when the proposed curriculum learning approach was applied.
@@ G.1 Convergence Speed across Model Size
G.2 Robustness on Noise Schedule
@@
For a more comprehensive ablation study, we also trained the diffusion model with different noise schedules. In contrast to cosine scheduling, the $\beta_{t}$ is set by uniformly dividing the interval $[0.0001,0.02]$, and the $C_{i}$ are obtained corresponding to SNR on a linear schedule. As shown in Table. A, our approach improves the performance with cosine and linear noise schedulers.
G.3 Qualitative Results from DiT-L/2 with 2M iterations
In Figures E-I, we present images generated by DiT-L using our curriculum training method for 2M iterations. The results demonstrate that our method produces highly realistic images.
Figure E: Uncurated 256$\times$256 DiT-L/2 samples.
Classifier-free guidanzce scale = 2.0.
Class label = “golden retriever" (207)
Figure F: Uncurated 256$\times$256 DiT-L/2 samples.
Classifier-free guidance scale = 2.0.
Class label = “panda" (388)
Figure G: Uncurated 256$\times$256 DiT-L/2 samples.
Classifier-free guidance scale = 2.0.
Class label = “cliff drop-off" (972)
Figure H: Uncurated 256$\times$256 DiT-L/2 samples.
Classifier-free guidance scale = 2.0.
Class label = “lake shore" (975)
Figure I: Uncurated 256$\times$256 DiT-L/2 samples.
Classifier-free guidance scale = 2.0.
Class label = “lion" (291)
Appendix H Broader Impacts
@@
Appendix I Limitations
In this work, we demonstrated the varying difficulties of denoising tasks through empirical results on various diffusion frameworks and proposed a curriculum learning approach that effectively enhances diffusion model training. While we have shown the robustness of our method’s hyperparameters in improving vanilla diffusion training, there is potential for further improvement. Specifically, curriculum learning methods that utilize smaller hyperparameters and adjust dynamically based on the model itself could yield better results. We acknowledge the validity of this direction and consider it a promising avenue for future work.